OpenStudy (anonymous):

solve the equation. 2(3.5+6)=2.5n-2

OpenStudy (anonymous):

Let's simplify first; 2(3.5+6)=2(9.5)=19. 19= 2.5n-2, add 2 to both sides, so that we get 21= 2.5n, then divide both by 2.5; 21/2.5 = 8.4 = 2.5n/2.5=n, n= 8.4